Need help with all the incorrect answers (I put wrong answers in Bold) please!! Answer for selling expense 2018 is not $111,000 either.

Problem 13-9

Pearl Company offers an MP3 download (seven-single medley) as a premium for every 5 candy bar wrappers presented by customers together with $2.90. The candy bars are sold by the company to distributors for 30 cents each. The purchase price of each download code to the company is $2.65. In addition, it costs 50 cents to distribute each code. The results of the premium plan for the years 2017 and 2018 are as follows. (All purchases and sales are for cash.)

2017

2018

MP3 codes purchased

300,000

396,000

Candy bars sold

2,803,800

2,967,500

Wrappers redeemed

1,440,000

1,800,000

2017 wrappers expected to be redeemed in 2018

348,000

2018 wrappers expected to be redeemed in 2019

420,000

Your answer is partially correct. Try again.

Prepare the journal entries that should be made in 2017 and 2018 to record the transactions related to the premium plan of the Pearl Company. (If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts. Credit account titles are automatically indented when amount is entered. Do not indent manually. Round answers to 0 decimal places, e.g. 1,525.)

Account Titles and Explanation

Debit

Credit

2017

Inventory of Premiums

795000

Cash

795000

(To record the premium inventory.)

Cash

841140

Sales Revenue

841140

(To record the sales.)

Cash

691200

Premium Expense

72000

Inventory of Premiums

763200

(To record the expense associated with the sale.)

Premium Expense

17400

Premium Liability

17400

(To record the premium liability.)

2018

Inventory of Premiums

1049400

Cash

1049400

(To record the premium inventory.)

Cash

890250

Sales Revenue

890250

(To record the sales.)

Cash

702600

Premium Expense

234000

Premium Liability

17400

Inventory of Premiums

954000

(To record the expense associated with the sale.)

Premium Expense

21000

Premium Liability

21000

(To record the premium liability.)

Your answer is partially correct. Try again.

Indicate the amounts for each accounts, and classifications of the items related to the premium plan that would appear on the balance sheet and the income statement at the end of 2017 and 2018.

Amount

Account

2017

2018

Classification

Inventory of Premiums

$31800

Entry field with correct answer

$95400

Entry field with incorrect answer

Entry field with correct answerCurrent Asset

Premium Liability

Entry field with correct answer17400

Entry field with correct answer21000

Entry field with correct answerCurrent Liability

Premium Expense

Entry field with correct answer89400

Entry field with incorrect answer255000

is not $111,000 either.

Entry field with correct answerSelling Expense
Answer #1
During 2017
Inventory (300,000*2.65) $795,000
Cash $795,000
Cash (2,803,800*.30) $841,140
Sales Revenue $841,140
Cash ((600,000-(240,000*.50)) $691,200
Premium Expense $72,000
Inventory of Premiums $763,200
Computation of premium expense (1440000/5*2.65) $763,200
     Postage (288,000*.50) $144,000
                                                   Total: $907,200
     Less: Cash Received (288,000*2.90) $835,200
     Premium Expense for MPS downloads issued $72,000
Premium Expense (348,000/5) *(2.65+.50-2.90) = 14,500 $17,400
Premium Liability $17,400
During 2018
Inventory of Premiums (396,000*2.65) $1,049,400
Cash $1,049,400
Cash (2,967,500*.30) $890,250
Sales Revenue $890,250
Cash $864,000
Premium Liability $17,400
Premium Expense $72,600
Inventory of Premiums $954,000
Computation of premium expense (1800000/5*2.65) $954,000
    Postage (360,000*.50) $180,000
$1,134,000
Less: Cash Received (1,800,000/5) *2.90 $1,044,000
Premium Expense of Codes issued $90,000
Less: Outstanding Claims at 12/31/2017 charged to 2017, but redeemed in 2018 $17,400
Premium Expense Chargeable to 2018 $72,600
Premium Expense (420,000/5) *(2.65+.50-2.90) $21,000
Premium Liability $21,000

ques 2

Amount
Account 2017 2018 Classification
Inventory Premiums $31,800.00 $127,200.00 Current Asset
Premium Liability $17,400.00 $21,000.00 Current Liability
Premium Expense $89,400.00 $93,600.00 Selling Expense
795000-763200=31800
17400+72000=89400
31800+1049400-954000=127200
21000+72600=93600